If you know that two legs of a right triangle are 10 and 11 units long, what is the triangle's area? If necessary, round your answer to two decimal places.
Answer by macston(5194)

In a right triangle, the legs are base and height and area=1/2bh
Area=1/2bh=1/2(10 units)(11 units)=1/2(110 sq units)=55 sq units
ANSWER The Area is 55 square units.